OpenStudy (anonymous):

Three sides of a triangle measure 10m, 12m, and 18m. Find the largest angle of the triangle to the nearest degree?

OpenStudy (zzr0ck3r):

For a triangle with side lengths \(a,b,c\) and angles \(A, B, C\), where a is the side opposite of A, b is the side opposite of B ..... If we know the three lengths, as you do, then we get the angles by these equations. \(A=\arccos\left(\frac{b^2+c^2-a^2}{2bc}\right) \\ B=\arccos\left(\frac{a^2+c^2-b^2}{2ac}\right) \\ C=180°-A-B\)